OpenStudy (anonymous):

To skate at Roller Heaven, each person must pay a membership fee (m) and a fee for each session (f). Art attended 5 sessions and paid $25 in all. Later, when he had attended 10 sessions, he paid $40 in all. How much is the membership fee and how much are the sessions. Show how to solve this, using any method.

OpenStudy (anonymous):

this question is sacreligious

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

ha jk "roller heaven"

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

25=m+5f 40=m+10f

OpenStudy (anonymous):

thx <3

OpenStudy (anonymous):

k